A large positively-charged object with charge q_+
Answer #1

From the diagram

Force between the charged objects is given as

F = k q1 q2/r2 =(9 x 109) (54.3 x 10-9) (4.25 x 10-6) / (0.225)2 = 0.041 N

Along Y-direction, force equation is given as

T Cos\theta = mg                              eq-1

Along X-direction, force equation is given as

T Sin\theta = F                              eq-2

Dividing eq-2 by eq-1

tan\theta = F/mg

tan\theta = 0.041 / (0.0165x 9.8)

\theta = 14.25